OpenStudy (anonymous):

Please help check answers! I'll give a medal! (x + 6)^2

OpenStudy (anonymous):

Would it be x36?

OpenStudy (anonymous):

x^2 +36

OpenStudy (anonymous):

(x+ 6)(x+6) Now use the Foil method.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

You can use the Foil method or know the identity a^ + 2ab + b^2

OpenStudy (anonymous):

\( (x+6)(x+6) = x^2 + 6x + 6x + 36 = x^2+12x+36\)

OpenStudy (anonymous):

That's it? Simplified?

OpenStudy (anonymous):

\( (x+6)(x+6) = x^2+12x+36\)

OpenStudy (anonymous):

Or we can say \( (x+6)^2 = x^2+12x+36\)

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Cool.

OpenStudy (anonymous):

Any other questions?

OpenStudy (anonymous):

No thank you you helped a lot :)